ON COMBINATORIAL DESCRIPTIONS OF HOMOTOPY GROUPS OF ΣK(π, 1)

Jie Wu

Abstract. We give a combinatorial description of homotopy groups of ΣK(π, 1). In particular, all of the homotopy groups of the 3-sphere are combinatorially given.

Typeset by AMS-TEX